OpenStudy (anonymous):

A boat running upstream takes 8 hours 48 minutes to cover a certain distance, while it takes 4 hours to cover the same distance running downstream. What is the ratio between the speed of the boat and speed of the water current respectively? [A]. 2 : 1 [B]. 3 : 2 [C]. 8 : 3 [D]. Cannot be determined [E]. None of these

OpenStudy (anonymous):

you can't solve this. the distance must be given.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Answer: Option C

OpenStudy (anonymous):

c is the answer?

OpenStudy (anonymous):

but how|? it confuses me.

OpenStudy (anonymous):

ya but dont kno steps

OpenStudy (anonymous):

Let the man's rate upstream be x kmph and that downstream be y kmph. Then, distance covered upstream in 8 hrs 48 min = Distance covered downstream in 4 hrs. 8 hrs. and 48 min is equal to 528/60 or 44/5 44x/5 = 4y y=11/5 x Required ratio = (y + x) /2 : (y - x)2 y=11/5x (11/5x+x)/2 : (11/5x - x)/2 16/5 x(1/2) : (6/5 x ) (1/2 8/5 x: 3/5 x cancel 5 so, the answer would be 8:3

OpenStudy (anonymous):

do you get it amit?

OpenStudy (anonymous):

STILL THERE?

OpenStudy (anonymous):

sry ... was busy in sm other wrk.. will hav a look thanks

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

Required ratio = (y + x) /2 : (y - x)2??

OpenStudy (anonymous):

think its ratio of downstream to upstream correct?

OpenStudy (anonymous):

yeah that is their ratio. I divided 2 in both sides.

OpenStudy (anonymous):

okies tq got it :)

OpenStudy (anonymous):

remember that if we divide a number to one side of the equation, we must also do the same way on the other side. it can't affect our answer.